Kyky232:

In ΔABC, ∠C measures 46° and the values of a and c are 10 and 9, respectively. Find the remaining measurements of the triangle, and round your answers to the nearest tenth. ∠A = 82.2°, ∠B = 62.8°, b = 17.1 ∠A = 53.1°, ∠B = 80.9°, b = 12.4 ∠A = 53.1°, ∠B = 80.9°, b = 17.1 ∠A = 82.2°, ∠B = 62.8°, b = 12.4 uwu plz help

Kyky232:

@AZ

snowflake0531:

do you know how to do the sin, cos, tan stuff?

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 do you know how to do the sin, cos, tan stuff? \(\color{#0cbb34}{\text{End of Quote}}\) No <3

snowflake0531:

ha, well, i can't explain that without confusing myself with the explanation lol AZ is on

Kyky232:

otay *patiently waits for az*

snowflake0531:

lollll fine, welllll, i'll say that sin(x) = opposite/hypotenuse cos(x)=adjacent/hypotenuse tan(x)=opposite/adjacent

AZ:



AZ:

We're not dealing with a 90 degree triangle so we have to use some other formulas. Have you learned law of sine/law of cosine

1 attachment
Kyky232:

yessir/ma'am/ fam I have learned that law :)

AZ:

So we know 2 sides and 1 angle so far We can use law of sine to find the angle opposite A \(\dfrac{\sin A}{a} = \dfrac{\sin C}{c}\) we know that a = 10 sin C = sin (46) c = 9 so can you solve for A?

Kyky232:

I think?

AZ:

Let's put the numbers in \(\dfrac{\sin A}{10} = \dfrac{\sin(46)}{9}\) Use a calculator and can you tell me what sin(46) =

Kyky232:

0.71933980033?

Kyky232:

well, rounding it would be .72

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Kyky232 0.71933980033? \(\color{#0cbb34}{\text{End of Quote}}\) let's use all the digits we can for now divide by 9 now

Kyky232:

0.07992664448

AZ:

now we have \(\dfrac{\sin A}{10} = 0.07992664448\) Multiply both sides by 10

Kyky232:

SinA = 0.7992664448

Kyky232:

Then we inverse Sine to get the angle of A

snowflake0531:

yeppi, correct

AZ:

Exactly! So what is arcsin(0.7992664448)

snowflake0531:

what the heck is arcsin.... don't you just do sin^-1 ???

snowflake0531:

oh, just inverse lol

Kyky232:

it would be 53.1 degrees

Kyky232:

I know what arson is :)

AZ:

sin^(-1) is the same as arcsin haha

snowflake0531:

nice

Kyky232:

nice

snowflake0531:

this is geometry, you're in 10th grade?

Kyky232:

8th :)

snowflake0531:

ha,, sameee

AZ:

so yeah 53.1 degrees for A is correct now we could first use law of cosines to find side B and then use law of sine to find angle b but we're smarter than that. We know all the angles of a triangle add up to 180 so 53.1 + b + 46 = 180 angle b = ?

Kyky232:

80.9 would be angle B

Kyky232:

am big brain B)

AZ:

now use law of sine again to calculate side B \(\dfrac{\sin(46)}{9} = \dfrac{\sin(80.9)}{B}\)

AZ:

now use law of sine again to calculate side B \(\dfrac{\sin(46)}{9} = \dfrac{\sin(80.9)}{B}\)

snowflake0531:

lol, when she doesn't reply-

Kyky232:

b = 12.4

AZ:

Good job!